OpenStudy (anonymous):

name 5 fractions that are near, but not equivalent to 4/12

OpenStudy (anonymous):

O_O

OpenStudy (anonymous):

there are many.for instance \[\frac{ 101 }{ 300 },\frac{ 103 }{ 300 },\frac{ 107 }{ 300 },\frac{ 100 }{ 301 },\frac{ 101 }{ 301 },...\] still closer \[\frac{ 1001 }{ 3000 },...\]

OpenStudy (anonymous):

well 4/12 is equal to 1/3 so it will be 2/3

OpenStudy (anonymous):

2/3 means 200/300 or 2000/3000 but which i wrote they are much closer.